# DiffuseDef: Improved Robustness to Adversarial Attacks

### v1

@@ 4.2 Implementation and Settings
We train two DiffuseDef variants using FreeLB++ and RSMI models as base models considering their robust adversarial defense capabilities. In the diffusion layer, only one transformer encoder layer (Vaswani et al., 2017) is used. The maximum noising timestep during training is set to 30 for AGNews and QNLI datasets, and 10 for IMDB dataset, while at inference time, we only apply 5 denoising steps for . We follow (Ho et al., 2020) to use a linear schedule from to . The diffusion layer is trained for 100 epochs, with the base classifier parameters frozen for efficiency. During the diffusion training stage, the same train-dev splits are used as in the adversarial training stage, thus ensuring no data leakage. At inference time, the number of ensembles is set to 10. Appendix C lists the hyper-parameters for each dataset.
@@ 5.1 Adversarial Robustness
| Dataset | PLM | Method | Clean% | AUA% | #Query | ||||
| TF | TB | BA | TF | TB | BA | ||||
| AGNews | BERT-base | Fine-Tuned | 94.4 | 10.2 | 25.4 | 27.1 | 348 | 372 | 379 |
| InfoBERT | 95.0 | 35.5 | 39.1 | 42.6 | 377 | 397 | 397 | ||
| FreeLB++ | 95.0 | 54.7 | 56.5 | 44.6 | 426 | 430 | 390 | ||
| EarlyRobust | 94.4 | 35.6 | 37.2 | 45.7 | 475 | 516 | 533 | ||
| RSMI | 94.3 | 52.6 | 56.7 | 55.4 | 680 | 737 | 687 | ||
| DiffuseDef-FreeLB++ (Ours) | 94.8 | 84.5 | 86.0 | 84.6 | 877 | 972 | 910 | ||
| DiffuseDef-RSMI (Ours) | 93.8 | 82.7 | 83.3 | 84.4 | 894 | 1029 | 930 | ||
| RoBERTa-base | Fine-Tuned | 94.9 | 34.1 | 36.9 | 43.6 | 372 | 396 | 410 | |
| InfoBERT | 95.5 | 40.2 | 45.2 | 48.6 | 392 | 421 | 430 | ||
| FreeLB++ | 95.4 | 57.5 | 62.9 | 55.9 | 444 | 467 | 447 | ||
| RSMI | 93.1 | 64.2 | 66.4 | 67.4 | 774 | 861 | 808 | ||
| DiffuseDef-RSMI (Ours) | 92.9 | 82.9 | 83.5 | 82.2 | 905 | 925 | 1047 | ||
| IMDB | BERT-base | Fine-Tuned | 93.3 | 7.7 | 8.3 | 10.5 | 540 | 534 | 378 |
| InfoBERT | 93.9 | 29.2 | 25.4 | 30.7 | 642 | 644 | 390 | ||
| FreeLB++ | 94.3 | 44.2 | 39.6 | 40.6 | 784 | 829 | 426 | ||
| EarlyRobust | 92.7 | 49.7 | 46.8 | 43.8 | 2267 | 2788 | 1841 | ||
| RSMI | 90.9 | 60.0 | 54.4 | 51.1 | 2840 | 3455 | 2070 | ||
| DiffuseDef-FreeLB++ (Ours) | 94.4 | 82.1 | 83.0 | 84.0 | 3174 | 4348 | 2842 | ||
| DiffuseDef-RSMI (Ours) | 90.2 | 80.9 | 79.8 | 79.8 | 3590 | 4748 | 2901 | ||
| RoBERTa-base | Fine-Tuned | 94.6 | 21.3 | 17.9 | 13.6 | 587 | 671 | 493 | |
| InfoBERT | 94.8 | 30.9 | 27.9 | 21.8 | 681 | 760 | 549 | ||
| FreeLB++ | 95.3 | 46.0 | 42.1 | 33.9 | 829 | 974 | 637 | ||
| RSMI | 92.7 | 77.9 | 74.3 | 70.6 | 3443 | 4342 | 2619 | ||
| DiffuseDef-RSMI (Ours) | 92.4 | 84.7 | 84.1 | 84.3 | 3673 | 4782 | 3007 | ||
@@
In Table 1, we compare the adversarial robustness of DiffuseDef with baselines and SOTA methods on AGNews and IMDB datasets trained with BERT and RoBERTa. DiffuseDef consistently outperforms all other methods on both datasets across both PLMs, exhibiting substantial improvements in accuracy under attack. After applying diffusion training, the AUA score for both FreeLB++ and RSMI models improves significantly, with an average increase of 30% AUA against the three attack methods. Note that despite the robust adversarial performance of the RSMI model, especially when trained with RoBERTa on the IMDB dataset, it still benefits from DiffuseDef. When comparing the clean accuracies to its base model (i.e. FreeLB++ and RSMI), DiffuseDef only shows a minor decline, between 0.2 and 0.7 accuracy score, which indicates that it can preserve the clean text performance while improving adversarial robustness. Moreover, models trained with DiffuseDef show a much smaller gap between clean accuracy and accuracy under attack, and such difference can be reduced to less than 10% AUA.
@@
Another benefit of DiffuseDef is the increased number of adversarial queries needed to obtain a successful attack. Models applying DiffuseDef require over twice the number of queries on both datasets compared to the other methods. This increase is even larger on the IMDB dataset due to the longer text length. For example, DiffuseDef model requires on average over 3000 queries to achieve a successful attack while FreeLB++ only needs 400 to 800 queries. The substantial increase suggests that even if the attackers manage to construct a successful adversarial attack, they need 2x to 3x more time to find the attack on DiffuseDef than other models, affirming the improved robustness from diffusion training. In addition, we observe that the number of queries for denoising-based methods (i.e. RSMI, DiffuseDef) is generally higher than adversarial training-based methods (i.e. InfoBERT, FreeLB++). This is because denoising-based methods transform the hidden representations of the adversarial texts into a non-deterministic representation. The introduction of randomness in hidden states results in uncertainty in model logits, thus increasing the difficulty finding vulnerable words to attack (Wang et al., 2023).
5.2 Ablation - NLI Task
To understand how each component contributes to DiffuseDef, we conduct an ablation analysis on the QNLI dataset (Table 2). Compared to the fine-tuning baseline, FreeLB++ increases the AUA score from 21.5 to 45.6, showing the benefit of adversarial training. After applying diffusion training (with inference timestep ), the score is further improved to 49.2, showing that diffusion training complements adversarial training. Finally, ensembling enhances adversarial performance and improves the score to 66.7, with the number of queries growing from 392 to 485. Similar improvements in both AUA and number of queries is found with the RSMI model after applying diffusion training and ensembling, which validates that the two components are complementary and that DiffuseDef is compatible with multiple SOTA defense methods.
| Method | Clean% | AUA% | #Query |
| Fine-Tuned (BERT) | 90.8 | 21.5 | 195 |
| FreeLB++ | 90.3 | 45.6 | 253 |
| + diffusion training | 90.2 | 49.2 | 392 |
| + ensembling | 90.3 | 66.7 | 485 |
| RSMI | 87.4 | 35.2 | 314 |
| + diffusion training | 86.5 | 40.0 | 353 |
| + ensembling | 86.4 | 55.5 | 459 |
5.3 Robustness w.r.t Token Length
Figure 2 provides comparison of defense rate for different models by token length on the IMDB dataset. The defense rate is calculated as the percentage of test examples in which TextFooler fails to construct a successful attack. All models except RSMI show a consistent trend that the defense rate declines as the texts lengthen. This trend can be attributed to the nature of adversarial attacks as longer texts allow for the generation of more adversarial examples. Specifically, adversarial training defense methods like InfoBERT and FreeLB++ show poor performance on longer texts (more than 300 tokens), with the defense rate reduced to near 0. This drastic decline indicates that given an adequate number of queries, the attacker is guaranteed to find a successful attack to fool these models. Similarly, EarlyRobust exhibits a performance drop on long texts as it is based on FreeLB training. RSMI, however, performs worse on short texts, but its defense rate increases as the text length grows. Compared to all SOTA defense approaches, the two DiffuseDef variants show a more steadily declining trend and maintain a higher defense rate across all token lengths, i.e. DiffuseDef is more robust to input text length.
@@ 5.4 Effect of Additional Denoising Steps
In Figure 3 we study how the inference denoising steps can affect the adversarial performance. For the DiffuseDef model without ensembling, both AUA score and the number of queries required to attack increase as the inference denoising step is larger. As the denoising step grows from 1 to 30, the AUA score improves from 58 to 65 while the number of attack queries grows from 430 to 780. In contrast, for DiffuseDef with ensembling, the model maintains a stable but robust performance in AUA and number of queries, regardless of the increase of . Considering that the ensembling introduces a notable performance increase, the DiffuseDef model is likely to be hitting an upperbound in both metrics, thus no further improvement is reached by increasing the denoising steps. However, it also shows that with ensembling, DiffuseDef can be applied with a smaller for better efficiency while maintaining a robust adversarial performance.
@@ 5.5 Ensembling Diffused Hidden Representations
In DiffuseDef the text hidden state is diffused and ensembled to form a denoised hidden representation, which contributes significantly to the improved adversarial robustness. In this section, we study how the ensembling diffused hidden representation helps defend against adversarial attacks.
As mentioned in Section 2.1, attack methods need to first rank token importance based on its influence on prediction. Specifically, the importance score is calculated by comparing the change of model prediction probablities after removing each word. In Figure 4, we compare the distribution of max token importance score between FreeLB++ and its DiffuseDef counterpart. Both FreeLB++ and DiffuseDef show a long-tail distribution with over 80 percent examples having a max token importance score below 0.1. This suggests that in most cases changing one single token will not significantly alter the prediction for both models. However, DiffuseDef shows a notably lower percentage of tokens when the max importance score is between 0.9 and 1, where the attacker can easily find the vulnerable token to construct adversarial examples. This difference shows that DiffuseDef can complicate the process of important word searching, which accounts for the increased number of queries required for a successful attack.
@@ 5.6 Efficiency of DiffuseDef
| Method | Params | FLOPS |
| Fine-Tuned (BERT) | 110M | 46G |
| EarlyRobust | 82M | 32G |
| FreeLB++ | 110M | 46G |
| InfoBERT | 110M | 46G |
| RSMI | 110M | 92G |
| RanMask () | 110M | 459G |
| SAFER ( | 110M | 459G |
| DiffuseDef () | 120M | 96G |
| DiffuseDef () | 120M | 267G |
Given that DiffuseDef adds additional denoising and ensembling steps during inference, it inevitably increases the computation time compared to its base model. To study its efficiency, we report the number of model parameters and inference FLOPS in Table 4. In addition to the defense methods in Table 1, we also compare the efficiency of DiffuseDef with two other SOTA ensembling-based defense methods, i.e. RanMask (Zeng et al., 2023) and SAFER Ye et al. (2020).
All SOTA models have the same number of parameters as the fine-tuned BERT model, except EarlyRobust which applies attention head pruning for better efficiency. DiffuseDef, with 1 additional diffusion layer, increases the number of parameters from 110M to 120M. DiffuseDef requires more inference FLOPS than non ensembling-based baselines such as FreeLB++ and EarlyRobust. With and , the FLOPS for DiffuseDef doubles from 46G to 96G, nevertheless, this number is close to RSMI model (92G FLOPS) as it requires gradient information during inference. Despite this increase, DiffuseDef is more efficient than ensembling-based methods like RanMask and SAFER which need to go through a full forward pass for all ensembles. With the same ensembling number of 10, both RanMask and SAFER require 459G FLOPS, which is 10x the number for BERT baseline. In contrast, even with increased to 5, DiffuseDef can be run faster with 267G FLOPS, showing that it can mitigate the efficiency problem from ensembling while maintaining the benefit of improved robustness.
@@ 6 Conclusions
We propose a novel adversarial defense method, DiffuseDef, which combines adversarial training, diffusion training, and ensembling to improve model robustness to adversarial attacks. DiffuseDef can build on any existing adversarial training method, training an additional diffusion layer to predict and remove randomly sampled noise at a given timestep. During inference, the diffusion layer is used to denoise the adversarial hidden states, which are ensembled to construct a robust text representation. Our experiments validate the effectiveness and efficiency of DiffuseDef, which significantly outperforms SOTA on three common adversarial attack methods. Analysis shows that DiffuseDef makes it difficult to find vulnerable tokens to attack, and also reduces the difference between the hidden representations of clean and adversarial texts.
7 Limitations
Scope
Our experiments focus on defending against three common black-box adversarial attack methods, while whether DiffuseDef improves model robustness against white-box attacks is unclear. White-box attacks have access to model parameters and can utilize gradient information to construct adversarial examples more efficiently than black-box attacks. Defending against white-box attacks is more challenging, and we consider this as a future direction of DiffuseDef.
Comparison with additional approaches
Due to the length limit, we do not compare against all current approaches. However we do compare with the SOTA methods with best adversarial robustness based on our preliminary experiments.
Efficiency
Despite the fact that DiffuseDef is more efficient than existing ensembling-based methods, it still requires more model parameters and inference FLOPS than non-ensembling-based models to achieve a better robustness. Future directions of this work might involve efforts to reduce the size of diffusion layer and number of ensembles to make DiffuseDef more efficient.
8 Ethical Considerations
In this paper we propose a new method DiffuseDef which uses a diffusion layer as a denoiser to provide robust and efficient text representation. We demonstrate that the proposed method could significantly improve the robustness of NLP systems to adversarial attacks. However, DiffuseDef cannot defend against all adversarial attacks without limitations (e.g. number of perturbed words, semantic similarity between original and adversarial examples). Potential risks might include creation of new adversarial attacks devised specifically for DiffuseDef.

### v2

@@ 4.2 Implementation and Settings
We train two DiffuseDef variants using FreeLB++ and RSMI models as base models considering their robust adversarial defense capabilities. In the diffusion layer, only one transformer encoder layer (Vaswani et al., 2017) is used. The maximum noising timestep during training is set to 30 for AGNews and QNLI datasets, and 10 for the IMDB dataset, while at inference time, we only apply 5 denoising steps for . We follow (Ho et al., 2020) to use a linear schedule from to . The base encoder and classifier are first fine-tuned for 10 epochs, and the diffusion layer is trained for 100 epochs, with the base encoder and classifier parameters frozen for efficiency. During the diffusion training stage, the same train-dev splits are used as in the adversarial training stage, thus ensuring no data leakage. At inference time, the number of ensembles is set to 10. Appendix C lists the training hyperparameters.
@@ 5.1 Adversarial Robustness
| Dataset | Method | Clean% | AUA% | #Query | ||||
| TF | TB | BA | TF | TB | BA | |||
| AGNews | Fine-Tuned | 94.4 | 10.2 | 25.4 | 27.1 | 348 | 372 | 379 |
| InfoBERT (Wang et al., 2021) | 95.0 | 35.5 | 39.1 | 42.6 | 377 | 397 | 397 | |
| FreeLB++ (Li et al., 2021) | 95.0 | 54.7 | 56.5 | 44.6 | 426 | 430 | 390 | |
| EarlyRobust (Xi et al., 2022) | 94.4 | 35.6 | 37.2 | 45.7 | 475 | 516 | 533 | |
| RMLM (Wang et al., 2023) | 94.3 | 45.6 | 44.4 | 54.1 | 642 | 738 | 705 | |
| ATINTER (Gupta et al., 2023) | 94.2 | 68.0 | 59.0 | 81.0 | 527 | 235 | 122 | |
| RSMI (Moon et al., 2023) | 94.3 | 52.6 | 56.7 | 55.4 | 680 | 737 | 687 | |
| DiffuseDef-FreeLB++ (Ours) | 94.8 | 84.5 | 86.0 | 84.6 | 877 | 972 | 910 | |
| DiffuseDef-RSMI (Ours) | 93.8 | 82.7 | 83.3 | 84.4 | 894 | 1029 | 930 | |
| IMDB | Fine-Tuned | 93.3 | 7.7 | 8.3 | 10.5 | 540 | 534 | 378 |
| InfoBERT (Wang et al., 2021) | 93.9 | 29.2 | 25.4 | 30.7 | 642 | 644 | 390 | |
| FreeLB++ (Li et al., 2021) | 94.3 | 44.2 | 39.6 | 40.6 | 784 | 829 | 426 | |
| EarlyRobust (Xi et al., 2022) | 92.7 | 49.7 | 46.8 | 43.8 | 2267 | 2788 | 1841 | |
| RMLM (Wang et al., 2023) | 93.6 | 48.8 | 47.9 | 47.1 | 2554 | 3258 | 2064 | |
| ATINTER (Gupta et al., 2023) | 94.3 | 25.0 | 15.0 | 39.0 | 1173 | 645 | 637 | |
| RSMI (Moon et al., 2023) | 90.9 | 60.0 | 54.4 | 51.1 | 2840 | 3455 | 2070 | |
| DiffuseDef-FreeLB++ (Ours) | 94.4 | 82.1 | 83.0 | 84.0 | 3174 | 4348 | 2842 | |
| DiffuseDef-RSMI (Ours) | 90.2 | 80.9 | 79.8 | 79.8 | 3590 | 4748 | 2901 | |
| QNLI | Fine-Tuned | 90.8 | 21.5 | 15.5 | 13.3 | 195 | 206 | 177 |
| InfoBERT (Wang et al., 2021) | 91.2 | 27.8 | 22.8 | 18.7 | 217 | 232 | 201 | |
| FreeLB++ (Li et al., 2021) | 90.3 | 45.6 | 40.2 | 30.5 | 253 | 279 | 226 | |
| EarlyRobust (Xi et al., 2022) | 89.2 | 24.3 | 21.2 | 19.1 | 265 | 292 | 255 | |
| ATINTER (Gupta et al., 2023) | 90.4 | 43.0 | 24.8 | 34.0 | 185 | 115 | 89 | |
| RSMI (Moon et al., 2023) | 87.4 | 35.2 | 30.9 | 28.2 | 314 | 353 | 314 | |
| DiffuseDef-FreeLB++ (Ours) | 90.3 | 66.7 | 65.3 | 64.4 | 485 | 587 | 543 | |
| DiffuseDef-RSMI (Ours) | 86.4 | 55.5 | 54.8 | 57.2 | 459 | 569 | 528 | |
@@
In Table 1, we compare the adversarial robustness of DiffuseDef with baselines and SOTA methods on AGNews, IMDB, and QNLI datasets trained with BERT (results for RoBERTa in Appendix E). DiffuseDef consistently outperforms all other methods on all three datasets across the three attacks, exhibiting substantial improvements in accuracy under attack. After applying diffusion training, the AUA score for both FreeLB++ and RSMI models improves significantly, with an average increase of 30% AUA against the three attack methods. When comparing the clean accuracies to its base model (i.e., FreeLB++ and RSMI), DiffuseDef only shows a minor decline, between 0.2 and 0.7 accuracy score, which indicates that it can preserve the clean text performance while improving adversarial robustness. Moreover, models trained with DiffuseDef show a much smaller gap between clean accuracy and accuracy under attack, and such a difference can be reduced to less than 10% AUA.
@@
Another benefit of DiffuseDef is the increased number of adversarial queries needed to obtain a successful attack. Models applying DiffuseDef require over twice the number of queries on all three datasets compared to the other methods. This increase is even larger on the IMDB dataset due to the longer text length. For example, the DiffuseDef model requires on average over 3000 queries to achieve a successful attack, while FreeLB++ can only defend for 400 to 800 queries. The substantial increase suggests that even if the attackers manage to construct a successful adversarial attack, they need 2x to 3x more time to find the attack on DiffuseDef than other models, affirming the improved robustness from diffusion training. In addition, we observe that the number of queries for denoising-based methods (i.e., RSMI, RMLM, and DiffuseDef) is generally higher than adversarial training-based methods (i.e., InfoBERT, FreeLB++). This is because denoising-based methods transform the hidden representations of the adversarial texts into a non-deterministic representation. The introduction of randomness in hidden states results in uncertainty in model logits, thus increasing the difficulty of finding vulnerable words to attack (Wang et al., 2023).
5.2 Robustness against White-box Attacks
In addition to black-box attacks, we also measure the robustness against white-box attacks, where the attackers can access the defender model’s gradients and parameters, which is a more challenging adversarial attack scenario than black-box attacks. In Table 2, we report AUA score under T-PGD (Yuan et al., 2023) and SemAttack Wang et al. (2022), the former being a variant of PGD attack (Madry et al., 2019) for text, while the latter adopts C&W attack (Carlini and Wagner, 2017) on embedding layers. We observe that white-box attacks result in lower AUA for all defense methods. For example, FreeLB++ performs poorly on IMDB when attacked by SemAttack, with the AUA score being 0.03. However, applying diffusion training still significantly improves adversarial robustness to white-box attacks. Note that even RSMI shows very robust performance against T-PGD444RSMI is incompatible with SemAttack because it requires gradients of the embedding layer, but SemAttack skips the embedding layer by sending embeddings directly to the target model., it can still further improve from DiffuseDef.
| Method | AGNews | IMDB | ||
| T-PGD | Sem | T-PGD | Sem | |
| Fine-Tuned | 8.8 | 41.5 | 3.0 | 1.3 |
| InfoBERT | 32.1 | 47.1 | 25.4 | 2.0 |
| FreeLB++ | 19.6 | 58.2 | 15.5 | 3.4 |
| RSMI | 79.6 | n/a | 43.3 | n/a |
| DiffuseDef-FreeLB | 59.4 | 68.1 | 50.3 | 28.2 |
| DiffuseDef-RSMI | 81.7 | n/a | 48.4 | n/a |
5.3 Ablation analysis
To understand how each component contributes to DiffuseDef, we conduct an ablation analysis on the AGNews dataset by disabling ensembling, diffusion denoising, and adversarial training one at a time. Table 3 shows the ablation results, and we find that removing either component (e.g., ensembling and diffusion denoising) at inference time leads to a sharp decrease in accuracy under attack, which indicates that both denoising and ensembling contribute significantly to the improved robustness. Nevertheless, removing adversarial training, though also introducing a degraded AUA score, is less detrimental than the other inference time components, showing that DiffuseDef is not dependent on a specific adversarial training method.
| Method | TF | TB | BA |
| DiffuseDef | 84.5 | 86.0 | 84.6 |
| w/o ensembling | 64.2 | 65.8 | 54.9 |
| w/o diffusion denoising | 57.1 | 58.4 | 47.7 |
| w/o adv training | 80.3 | 80.9 | 80.5 |
5.4 DiffuseDef + Adversarial Augmentation
In Figure 2, we combine DiffuseDef with adversarial data augmentation and evaluate the robustness on AGNews. We first generate adversarial augmentation data using TextFooler and adversarially train the classifier with augmented training data. The augmented classifier is further performed with diffusion training, resulting in a DiffuseDef variant. We notice that adversarial training on augmented data helps improve the performance on all three attacks. After applying diffusion training, the adversarial performance significantly improves over the augmented model, which proves the flexibility of DiffuseDef to be applied with different types of adversarial training methods.
@@ 5.5 Effect of Additional Denoising Steps
In Figure 3, we study how the inference denoising steps can affect the adversarial performance. For the DiffuseDef model without ensembling, both the AUA score and the number of queries required to attack increase as the inference denoising step is larger. As the denoising step grows from 1 to 30, the AUA score improves from 58 to 65 while the number of attack queries grows from 430 to 780. In contrast, for DiffuseDef with ensembling, the model maintains a stable but robust performance in AUA and number of queries, regardless of the increase of . Considering that the ensembling introduces a notable performance increase, the DiffuseDef model is likely to be hitting an upper bound in both metrics, thus no further improvement is reached by increasing the denoising steps. However, it also shows that with ensembling, DiffuseDef can be applied with a smaller for better efficiency while maintaining a robust adversarial performance.
@@ 5.6 Ensembling Diffused Hidden Representations
In DiffuseDef, the text hidden state is diffused and ensembled to form a denoised hidden representation, which contributes significantly to the improved adversarial robustness. In this section, we study how the ensembling diffused hidden representation helps defend against adversarial attacks.
As mentioned in Section 2.1, attack methods need to first rank token importance based on its influence on prediction. Specifically, the importance score is calculated by comparing the changes in model prediction probabilities after removing each word. In Figure 4, we compare the density distribution of the max token importance score between FreeLB++ and its DiffuseDef counterpart. DiffuseDef shows a notably lower percentage than FreeLB++ when the max importance score is high, where the attacker can easily find the vulnerable token to construct adversarial examples. This difference shows that DiffuseDef can complicate the process of important word searching, which accounts for the increased number of queries required for a successful attack.
@@ 5.7 Efficiency of DiffuseDef
| Method | Params | FLOPS | Train |
| Fine-Tuned (BERT) | 110M | 46G | 1x |
| EarlyRobust | 82M | 32G | 0.8x |
| FreeLB++ | 110M | 46G | 10.5x |
| InfoBERT | 110M | 46G | 6.5x |
| RSMI | 110M | 92G | 1.25x |
| RanMask () | 110M | 459G | 1.2x |
| SAFER ( | 110M | 459G | 1x |
| DiffuseDef () | 120M | 96G | 1.1x |
| DiffuseDef () | 120M | 267G | 1.1x |
Given that DiffuseDef adds additional denoising and ensembling steps during inference, it inevitably increases the computation time compared to its base model. To study its efficiency, we report the number of model parameters, inference FLOPS, and training time in Table 5. In addition to the defense methods in Table 1, we also compare the efficiency of DiffuseDef with two other SOTA ensembling-based defense methods, i.e., RanMask (Zeng et al., 2023) and SAFER Ye et al. (2020).
All SOTA models have the same number of parameters as the fine-tuned BERT model, except EarlyRobust, which applies attention head pruning for better efficiency. DiffuseDef requires more inference FLOPS than non-ensembling-based baselines such as FreeLB++ and EarlyRobust, due to the diffusion steps. With and , the FLOPS for DiffuseDef doubles from 46G to 96G, nevertheless, this number is close to the RSMI model as it requires gradient information during inference. Despite this increase, DiffuseDef is more efficient than ensembling-based methods like RanMask and SAFER, which require a full forward pass for all ensembles. With the same ensembling number of 10, both RanMask and SAFER require 459G FLOPS, which is 10x the number for the BERT baseline. In contrast, even with increased to 5, DiffuseDef can be run faster with 267G FLOPS, mitigating the efficiency problem from ensembling while maintaining the benefit of improved robustness.
Regarding training time, although training the diffusion layer requires more epochs, it costs similar time as training the classifier due to the small size of the diffusion layer. By contrast, adversarial training methods such as FreeLB++ and InfoBERT require much longer training time due to increased iterations of back-propagation per training batch.
@@ 6 Conclusions
We propose a novel adversarial defense method, DiffuseDef, which combines adversarial training, diffusion training, and ensembling to improve model robustness to adversarial attacks. DiffuseDef can build on any existing adversarial training method, training an additional diffusion layer to predict and remove randomly sampled noise at a given timestep. During inference, the diffusion layer is used to denoise the adversarial hidden states, which are ensembled to construct a robust text representation. Our experiments validate the effectiveness and efficiency of DiffuseDef, which significantly outperforms SOTA on common black-box and white-box adversarial attack methods, and maintains the performance on clean texts. Analysis shows that DiffuseDef makes it difficult to find vulnerable tokens to attack, and also reduces the difference between the hidden representations of clean and adversarial texts.
Limitations
Despite the fact that DiffuseDef is more efficient than existing ensembling-based methods, it still requires more model parameters and inference FLOPS than non-ensembling-based models to achieve better robustness. Future directions of this work might involve efforts to reduce the size of the diffusion layer and the number of ensembles to make DiffuseDef more efficient.
Ethical Considerations
In this paper, we propose a new method DiffuseDef, which uses a diffusion layer as a denoiser to provide robust and efficient text representation. We demonstrate that the proposed method could significantly improve the robustness of NLP systems to adversarial attacks. However, DiffuseDef cannot defend against all adversarial attacks without limitations (e.g., number of perturbed words, semantic similarity between original and adversarial examples). Potential risks might include the creation of new adversarial attacks devised specifically for DiffuseDef.